# OpenGraph: Towards Open Graph Foundation Models

### v2

@@ Abstract.
Graph learning has become indispensable for interpreting and harnessing relational data in diverse fields, ranging from recommendation systems to social network analysis. In this context, a variety of graph neural networks (GNNs) have emerged as promising methodologies for encoding the structural information of graphs. By effectively capturing the graph’s underlying structure, these GNNs have shown great potential in enhancing performance in graph learning tasks, such as link prediction and node classification. However, despite their successes, a significant challenge persists: these advanced methods often face difficulties in generalizing to unseen graph data that significantly differs from the training instances. In this work, our aim is to advance the graph learning paradigm by developing a general graph foundation model. This model is designed to understand the complex topological patterns present in diverse graph data, enabling it to excel in zero-shot graph learning tasks across different downstream datasets. To achieve this goal, we address several key technical challenges in our OpenGraph model. Firstly, we propose a unified graph tokenizer to adapt our graph model to generalize well on unseen graph data, even when the underlying graph properties differ significantly from those encountered during training. Secondly, we develop a scalable graph transformer as the foundational encoder, which effectively captures node-wise dependencies within the global topological context. Thirdly, we introduce a data augmentation mechanism enhanced by a large language model (LLM) to alleviate the limitations of data scarcity in real-world scenarios. Extensive experiments validate the effectiveness of our framework. By adapting our OpenGraph to new graph characteristics and comprehending the nuances of diverse graphs, our approach achieves remarkable zero-shot graph learning performance across various settings and domains.
* Chao Huang is the Corresponding Author.
copyright: none
1. Introduction
@@
Graph learning has established itself as a core and indispensable methodology for unlocking the inherent power of relational data across a wide spectrum of fields. Its versatility and applicability span diverse domains, including recommender systems (he2020lightgcn, ), social network analysis (sankar2021graph, ), citation networks (lv2021we, ), and transportation networks (wang2020traffic, ). To effectively capture the intricate structures inherent in graphs, the adoption of Graph Neural Networks (GNNs) has become pervasive in learning representations for graph-structured data. GNNs are built on the foundation of recursive message passing, leveraging the inter-dependencies of nodes to incorporate high-order connectivities into the learned graph representations (ying2018graph, ; jin2020graph, ).
One of the primary challenges faced by current end-to-end graph neural networks is their heavy reliance on labeled data. This strong dependence on labeled data presents a primary obstacle in real-world situations, where obtaining a sufficient amount of accurately labeled data can be both expensive and challenging (liu2022graph, ; jin2021automated, ). Consequently, the labeled data that is available often exhibits problems of being sparse and of low quality. To tackle this issue, self-supervised learning (SSL) has emerged as a potential solution to enhance graph neural networks by leveraging augmented self-supervision signals. These methods typically adopt a learning pipeline that involves "pre-training" and "fine-tuning" with self-supervised pre-training tasks. In the context of self-supervised learning and pre-training on graphs, contrasive SSL has been the dominant approaches. For example, contrastive objectives are incorporated as self-supervised alignment loss to supplement the classical supervised task, Notable examples include DGI (velivckovic2018deep, ) and GraphCL (you2020graph, ). Further research, such as JOAO (you2021graph, ) and GCA (zhu2021graph, ), aims to automate the contrastive learning process through adaptive augmentation.
While graph pre-training techniques have demonstrated effectiveness in capturing intrinsic graph properties, they still encounter difficulties in achieving strong generalization from the pre-training phase to the downstream task-specific context (sun2023all, ), especially when confronted with distribution shifts in graph characteristics across diverse downstream domains (wu2021handling, ; gui2022good, ). For example, in recommender systems, there is a need to handle previously unseen user interaction graphs that arise from newly generated items, such as news articles or videos, in cold-start recommendation scenarios (chen2022generative, ). Furthermore, there is a desire to transfer the acquired knowledge from pre-trained graph domains to other downstream domains (zhang2022few, ). However, when these models are applied to unseen graphs, their performance significantly deteriorates. This decline can be attributed to the variations in graph characteristics observed in different downstream scenarios, including changes in node sets and feature semantics.
@@
Recent research has explored the use of prompt-tuning as a task-specific alternative to fine-tuning, aiming to bridge the gap between pre-training and downstream objectives (sun2022gppt, ; liu2023graphprompt, ; fang2022universal, ). By merging the pre-training and prompt-tuning frameworks, these approaches enable a more efficient alignment of the pre-trained model’s understanding with the specific requirements of the downstream task. However, it is important to note that most of these frameworks assume that the training and testing graph data share the same node set and feature semantics. In practical scenarios, variations in node sets and feature semantics are common across diverse downstream graph domains. Therefore, further exploration is needed to enhance the generalization power of graph models and ensure their adaptability to various real-world downstream graphs.
Considering the aforementioned motivations and challenges, the objective of this work is to push the boundaries of graph learning by developing a versatile and scalable graph model. The primary aim is to equip the model with the capability of zero-shot learning, enabling it to effectively learn from and make accurate predictions on previously unseen graphs. However, building such a graph model is not a trival task and requires addressing several key challenges.
  • C1: Node Token Set Shift across Different Graphs. One notable challenge in zero-shot graph learning is the shift in node token sets across different graphs. This discrepancy necessitates the model to proficiently reconcile the variations in node characteristics. It is crucial to generate universal graph tokens that possess the adaptability to effectively represent and comprehend arbitrary unseen graphs with diverse graph topological contexts.
  • C2: Efficient Node-wise Dependency Modeling. In the realm of graph learning, nodes in large-scale graphs exhibit intricate dependencies on each other. It is crucial to comprehend both local and global inter-dependencies among all nodes for accurate prediction and effective learning. Yet, when it comes to modeling universal dependencies over extensive graphs, the efficiency of node-wise dependency encoding becomes vital in enhancing the performance and scalability of graph models.
@@
  • C3: Domain-Specific Data Scarcity. Data scarcity is a widespread challenge across various downstream domain tasks, and it arises from multiple factors. Privacy concerns, for example, restrict the collection of domain-specific user behavior graphs, resulting in limited data availability. Consequently, it becomes crucial to develop label-less learning frameworks within graph foundation models to effectively understand the relevant context of downstream tasks under conditions of data scarcity.
Present Work. To overcome these challenges, we present OpenGraph, a flexible graph model that excels in zero-shot graph learning, allowing it to capture the universal and transferable structural patterns across multiple domains or datasets. To address the first challenge C1, we introduce a graph tokenizer that efficiently transforms input graphs into unified token sequences. This is achieved by incorporating a topology-aware projection scheme, enabling the generation of universal graph tokens from arbitrary graphs, regardless of variations in node token sets. To conquer the second challenge, C2, we have developed a highly scalable graph transformer equipped with efficient self-attention using anchor sampling. This innovative approach ensures computational efficiency through a two-stage self-attention process. Additionally, we employ token sequence sampling to optimize the training process, leveraging sampled token sequences from the current batch. This reduces the sequence length while preserving the crucial graph context.
@@
To tackle the challgne of C3, we introduce a novel approach that combines the power of large language models (LLMs) with data augmentation techniques for synthetic graph generation. Our goal is to enhance the pre-training process of our OpenGraph model by incorporating more downstream task-relevant context and gaining a deeper understanding of real-world graph characteristics. This is achieved through the generation of augmented graphs that closely resemble real-world instances. To achieve this, we employ an integration of the tree-of-prompt regularized with Gibbs sampling.
We conducted extensive experiments on a diverse range of publicly available graph datasets, encompassing various scales and types. The outcomes were truly remarkable, as our OpenGraph consistently showcased exceptional generalization capabilities across a wide array of settings. Particularly exciting is the fact that the zero-shot learning ability of OpenGraph surpassed baselines even in few-shot learning scenarios. This outcome lays the foundation for the development of graph foundation models, which have the potential to generalize effectively across diverse graph domains.
2. Preliminaries
Graph Representation Learning. A graph $\mathcal{G}$ consists of a set of nodes $\mathcal{V}$ and a set of edges $\mathcal{E}$. The edges $e=(v_{s},v_{t})\in\mathcal{E}$ represent observed relations between a source node $v_{s}$ and a target node $v_{t}$. Additionally, we define a feature matrix $\textbf{F}\in\mathbb{R}^{|\mathcal{V}|\times f}$ to capture the $f$-dimensional feature attributes of all nodes. Graph representation learning aims to capture meaningful node representations by transforming both the structural and attribute information into low-dimensional vector embeddings. These embeddings encode the patterns of inter-node dependencies and relationships, facilitating graph learning tasks such as link prediction and node classification.
The link prediction focuses on predicting the probability of establishing connections between node pairs $(v_{s},v_{t})$ that are not observed in the given data. The node classification task aims to assign a category to each node based on its features and relationships with other nodes. These tasks can be formally expressed:
$\mathop{\arg\min}_{\mathbf{\Theta}_{f}}\mathcal{L}_{\text{node}}=-\sum_{v_{s},y_{s}}\Big{(}{f(v_{s},y_{s})}\leavevmode\nobreak\ /{\sum_{y^{\prime}_{s}\neq y_{s}}f(v_{s},y^{\prime}_{s})}\Big{)},$
(1) $\mathop{\arg\min}_{\mathbf{\Theta}_{f}}\mathcal{L}_{\text{link}}=-\sum_{v_{s},v_{t}}e_{s,t}f(v_{s},v_{t})-(1-e_{s,t})f(v_{s},v_{t})$
@@ 2. Preliminaries
Here, The link labels $e_{s,t}\in\{0,1\}$ indicate if node $v_{s}$ is connected to node $v_{t}$. $f$ denotes the prediction model with a learnable parameter set $\mathbf{\Theta}_{f}$. The node labels $y_{s}\in\mathcal{C}$ indicate the groundtruth category for node $v_{s}$, while $\mathcal{C}$ denotes the set of candidate categories.
Zero-shot Graph Learning. Although current graph learning methods demonstrate promising results in standard graph learning tasks, they face limitations when it comes to effectively generalizing across diverse domains. When these models are applied to new and unseen graphs, their performance noticeably decreases. This decline can be attributed to the variations in graph characteristics that exist across different downstream scenarios, such as changes in the node set and feature semantics. Existing models heavily depend on parameters that are specifically tailored to these graph-specific tokens, which include feature transformations and node embeddings. As a result, these models encounter challenges in their ability to effectively generalize and adapt to new graphs that possess distinct graph tokens and relational semantics.
Recognizing the limitations of current approaches, our research focuses on zero-shot graph learning. This task involves training a model on a set of graphs and evaluating its performance on different test graphs that do not share any graph tokens. The goal is to assess the ability of graph models to learn generalized graph topological structures and node-wise dependencies. To provide a formal definition of the zero-shot graph learning task:
$\mathop{\arg\min}\nolimits_{f}\epsilon(\{\mathcal{G}_{t}\},f),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ \mathbf{\Theta}_{f}=\mathop{\arg\min}\nolimits_{\mathbf{\Theta}_{f}}\mathcal{L}(\{\mathcal{G}_{s}\},f),$
(2) $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{V}_{t}\cap\mathcal{V}_{s}=\varnothing,\leavevmode\nobreak\ \mathcal{E}_{t}\cap\mathcal{E}_{s}=\varnothing,\leavevmode\nobreak\ \mathbb{R}^{f_{t}}\neq\mathbb{R}^{f_{s}},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\mathcal{G}_{s},\mathcal{G}_{t}$
@@
The objective of this task is to develop a graph model, denoted as $f(\cdot)$, that effectively understands general and universal topological patterns. It should also discern contextually relevant patterns crucial for high performance in diverse downstream graph learning tasks. Notably, the training graphs ($\mathcal{G}_{s}$) and test graphs ($\mathcal{G}_{t}$) have no common nodes ($v\in\mathcal{V}$), edges ($e\in\mathcal{E}$), or node feature attribute semantics. This presents a unique challenge for the graph model to handle the significant distribution shift that occurs across different graph domains with entirely distinct datasets.
3. Methodology
In this section, we delve into the design details of the proposed OpenGraph framework. To offer a comprehensive overview of the model, Figure 1 illustrates the overall structure and components.
Figure 1. Overall architecture of the OpenGraph. i) A unified graph tokenizer efficiently converts any input graph into universal graph tokens, accommodating variations in node token sets. ii) The scalable graph transformer effectively captures global node-wise dependencies using anchor sampling and efficient self-attention mechanisms. iii) LLM-enhanced data augmentation techniques are employed for synthetic graph generation, addressing the issue of limited domain-specific data availability.
3.1. Unified Graph Tokenizer
@@
To overcome the challenges presented by diverse graphs that differ significantly in terms of nodes, edges, and feature semantics, our primary goal is to develop a graph tokenizer that effectively transforms input graphs into unified token sequences. In our graph tokenizer, each token represents a node and is accompanied by a semantic vector denoted as $\mathcal{G}\rightarrow{\textbf{e}_{0},\textbf{e}_{1},\cdots,\textbf{e}_{|\mathcal{V}|}}$. By utilizing a shared node representation space and employing a flexible sequence data structure, we aim to standardize the distribution of nodes across different graphs. Additionally, we merge the edge information into the unified node representations, disregarding differences between graphs. To address the variations in node features, we introduce a topological representation space that is generated through a consistent mapping procedure.
3.1.1. Smoothed High-Order Adjacency Matrix
In our OpenGraph model, we employ a unified graph tokenizer that utilizes the smoothed adjacency matrix, denoted as $\tilde{\textbf{A}}$, and a topology-aware projection function $\phi:\mathbb{R}^{|\mathcal{V}|}\rightarrow\mathbb{R}^{d}$. This tokenizer is responsible for handling the specific graph representation. To achieve this, we start with the original adjacency matrix $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ constructed from the edges $\mathcal{E}$. In our OpenGraph paradigm, the smoothing procedure for the adjacency matrix is carried out as follows:
(3) $\tilde{\textbf{A}}=\bar{\textbf{A}}^{1}+\bar{\textbf{A}}^{2}+\cdots\bar{\textbf{A}}^{L},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{\textbf{A}}=\textbf{D}^{-1/2}\textbf{A}\textbf{D}^{-1/2}$
@@ 3.1.1. Smoothed High-Order Adjacency Matrix
Here, we utilize the diagonal degree matrix D derived from the adjacency matrix A. It records the node degree of $v_{i}$ with $d_{i,i}$, and $d_{i,j}=0$ for $i\neq j$. With matrix D, we employ the Laplacian normalization $\bar{\textbf{A}}$ of A to ensure numerical stability. To capture high-order connectivity and address sparsely-observed node-wise relations, our OpenGraph combines the power of $\bar{\textbf{A}}$ at different orders. This allows us to obtain topology information $\tilde{\textbf{A}}$ for further processing, with $L$ representing the maximum power order considered.
3.1.2. Topology-aware Projection with Arbitary Graphs

### v4

@@ Abstract
Graph learning has become essential in various domains, including recommendation systems and social network analysis. Graph Neural Networks (GNNs) have emerged as promising techniques for encoding structural information and improving performance in tasks like link prediction and node classification. However, a key challenge remains: the difficulty of generalizing to unseen graph data with different properties. In this work, we propose a novel graph foundation model, called OpenGraph, to address this challenge. Our approach tackles several technical obstacles. Firstly, we enhance data augmentation using a large language model (LLM) to overcome data scarcity in real-world scenarios. Secondly, we introduce a unified graph tokenizer that enables the model to generalize effectively to diverse graph data, even when encountering unseen properties during training. Thirdly, our developed scalable graph transformer captures node-wise dependencies within the global topological context. Extensive experiments validate the effectiveness of our framework. By adapting OpenGraph to new graph characteristics and comprehending diverse graphs, our approach achieves remarkable zero-shot graph learning performance across various settings. We release the model implementation at https://github.com/HKUDS/OpenGraph.
OpenGraph: Towards Open Graph Foundation Models
Lianghao Xia University of Hong Kong aka_xia@foxmail.com                        Ben Kao University of Hong Kong kao@cs.hku.hk                        Chao Huangthanks: Chao Huang is the corresponding author. University of Hong Kong chuang7@hku.hk
1 Introduction
@@
Graph learning is a crucial methodology in various fields, such as recommender systems He et al. (2020), social network analysis Sankar et al. (2021), citation networks Lv et al. (2021), and transportation networks Wang et al. (2020). By utilizing Graph Neural Networks (GNNs) and recursive message passing, we capture the complex structures of graphs effectively. GNNs leverage inter-dependencies among nodes to incorporate high-order connectivities into learned graph representations Ying et al. (2018); Jin et al. (2020).
A primary challenge in current end-to-end graph neural networks is their heavy reliance on scarce and low-quality labeled data Liu et al. (2022); Jin et al. (2022). To overcome this, self-supervised learning (SSL) has emerged as a solution by leveraging augmented self-supervision signals. Contrastive SSL, exemplified by DGI Veličković et al. (2018b) and GraphCL You et al. (2020), incorporates contrastive objectives as self-supervised alignment loss. Recent advancements like JOAO You et al. (2021) and GCA Zhu et al. (2021) automate contrastive learning through adaptive augmentation. By integrating SSL techniques, we enhance graph neural networks with limited labeled data.
Graph pre-training excels at capturing intrinsic graph properties but struggle to effectively generalize to diverse downstream domains, particularly when faced with distribution shifts Sun et al. (2023); Wu et al. (2021c); Gui et al. (2022). For example, in recommender systems, handling previously unseen user interaction graphs in cold-start recommendation scenarios is crucial Chen et al. (2022a). Transferring knowledge from pre-trained graph domains to other downstream domains is desirable Zhang et al. (2022a). However, applying these models to unseen graphs results in significant performance deterioration due to variations in node sets and relation semantics across different scenarios.
@@
Recent research explores prompt-tuning as a task-specific alternative to fine-tuning, bridging the gap between pre-training and downstream objectives Sun et al. (2022); Liu et al. (2023); Fang et al. (2023). These approaches align the pre-trained model’s understanding with specific task requirements. However, practical scenarios involve variations in node sets and feature semantics across diverse downstream graph domains. Further exploration is needed to enhance graph models’ generalization and adaptability to real-world graphs.
This work aims to develop a scalable graph model that enables zero-shot learning, effectively making accurate predictions on unseen graphs. Building such a model poses significant challenges.
@@
  • C1: Domain-Specific Data Scarcity. Data scarcity is a common challenge across downstream domain tasks, driven by factors like privacy concerns. Limited availability of domain-specific user behavior graphs restricts data collection. Therefore, developing label-less learning frameworks within graph models is crucial to effectively understand the context of downstream tasks in the face of data scarcity.
  • C2: Node Token Set Shift. A key challenge in zero-shot graph learning is the shift in node token sets across graphs. This requires the model to reconcile variations in node characteristics. Generating universal graph tokens is crucial to effectively represent and comprehend diverse unseen graphs with different topological contexts.
  • C3: Efficient Graph Dependency Modeling. Nodes in large-scale graphs have complex dependencies. Understanding local and global inter-dependencies among all nodes is crucial for accurate prediction. Efficient node-wise dependency encoding is vital to enhance the performance and scalability of graph models.
@@
Present Work. To overcome the challenges, we introduce a graph model for zero-shot learning that captures universal and transferable structural patterns across multiple domains. To address C1, we propose combining large language models (LLMs) with data augmentation techniques for synthetic graph generation. By generating augmented graphs resembling real-world instances, we enhance the pre-training process of OpenGraph and gain a deeper understanding of downstream task contexts. This is achieved through the integration of tree-of-prompt regularization with Gibbs sampling.
To address C2, we propose a topology-aware graph tokenizer that generates universal graph tokens from arbitrary graphs. For C3, we develop a scalable graph transformer with efficient self-attention using anchor sampling. Our approach ensures computational efficiency through a two-stage self-attention process and optimizes training by leveraging token sequence sampling, reducing sequence length while preserving crucial context. We conducted extensive experiments on diverse datasets, demonstrating the remarkable generalization abilities of our model across various settings.
2 Preliminaries
Graph Learning. A graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\textbf{F})$ consists of a node set $\mathcal{V}={v_{i}}$, an edge set $\mathcal{E}={(v_{s},v_{t})}$, and node attributes $\textbf{F}\in\mathbb{R}^{|\mathcal{V}|\times f}$. Graph learning aims to produce node representations that encode both structural and attribute information. These embeddings are used for tasks such as link prediction and node classification, involving the prediction of node connections and categories, respectively. The corresponding losses to be minimized are:
$\mathcal{L}_{\text{link}}=\sum_{v_{s},v_{t}}(1-e_{s,t})f(v_{s},v_{t})-e_{s,t}f(v_{s},v_{t}),$
$\mathcal{L}_{\text{node}}=-\sum_{v_{s},y_{s}}\Big{(}{f(v_{s},y_{s})}\leavevmode\nobreak\ /{\sum_{y^{\prime}_{s}\neq y_{s}}f(v_{s},y^{\prime}_{s})}\Big{)}$ (1)
@@ 2 Preliminaries
where $e_{s,t}\in\{0,1\}$ denotes the link label for nodes $v_{s}$ and $v_{t}$, and $y_{s}\in\mathcal{C}$ indicate the groundtruth category for node $v_{s}$. Function $f$ denotes the prediction model with learnable parameters $\mathbf{\Theta}_{f}$.
Zero-shot Graph Learning. Current graph models excel in standard tasks but struggle to generalize across diverse domains. Their performance deteriorates when applied to new graphs with varying characteristics, such as node sets and features. To address these limitations, we focus on zero-shot graph learning, where a model is trained on a set of graphs and evaluated on different test graphs without shared graph tokens. It aims to assess the model’s ability to learn generalized topological structures and node-wise dependencies. Formally, we seek to minimize the error measurement $\epsilon({\mathcal{G}_{t}},f)$, with $\mathop{\arg\min}\nolimits_{f}$ denoting the optimization.
$\mathbf{\Theta}_{f}=$ $\mathop{\arg\min}\nolimits_{\mathbf{\Theta}_{f}}\mathcal{L}(\{\mathcal{G}_{s}\},f),$
$\mathcal{V}_{t}\cap\mathcal{V}_{s}=$ $\varnothing,\leavevmode\nobreak\ \mathcal{E}_{t}\cap\mathcal{E}_{s}=\varnothing,\leavevmode\nobreak\ \mathbb{R}^{f_{t}}\neq\mathbb{R}^{f_{s}}$ (2)
@@
This objective is to develop a universal graph modeling architecture $f(\cdot)$. Its parameters $\mathbf{\Theta}_{f}$ are learned by minimizng graph learning losses $\mathcal{L}$ on the training graphs $\{\mathcal{G}_{s}\}$. Notably, the training graphs and test graphs $\mathcal{G}_{t}$ have no common nodes, edges, or node features. This presents a unique challenge for the graph model to handle the significant distribution shift that occurs across different graph domains with entirely distinct datasets.
3 Methodology
This section presents the design details of the proposed OpenGraph framework. Figure 1 gives an overall illustration for the model. In appendix, we elaborate on how OpenGraph handles node classifiction (A.1.1) and graph features (A.1.2), detailed configurations of our scalable graph transformer (A.1.3), as well as the generation algorithms (A.2).
Figure 1: Overall model architecture of the OpenGraph framework.
3.1 Unified Graph Tokenizer
@@
To handle diverse graphs with varying nodes and features, our goal is to develop a graph tokenizer that transforms input graphs into unified token sequences: $\mathcal{G}\rightarrow\{\textbf{e}_{i}\}$. Each token represents a node accompanied by a semantic vector $\textbf{e}_{i}$. By utilizing a shared representation space and a flexible sequence structure, we aim to standardize distributions across graphs. Specifically, our tokenizer uses the smoothed adjacency matrix $\tilde{\textbf{A}}$, and a topology-aware projection function $\phi:\mathbb{R}^{|\mathcal{V}|}\rightarrow\mathbb{R}^{d}$.
3.1.1 Smoothed High-Order Adjacency
We start with the original adjacency matrix $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ built from edges $\mathcal{E}$. The smoothing procedure for the adjacency matrix is as follows:
$\tilde{\textbf{A}}=\bar{\textbf{A}}^{1}+\bar{\textbf{A}}^{2}+\cdots\bar{\textbf{A}}^{L},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{\textbf{A}}=\textbf{D}^{-\frac{1}{2}}\textbf{A}\textbf{D}^{-\frac{1}{2}}$ (3)
@@ 3.1.1 Smoothed High-Order Adjacency
For numerical stability, we use Laplacian normalization $\bar{\textbf{A}}$ with the diagonal degree matrix D of adjacency A. To capture high-order connectivity and sparse node-wise relations, OpenGraph combines $\bar{\textbf{A}}$ at different orders. This provides us with topology information for further processing, with $L$ representing the maximum power order considered.
3.1.2 Topology-aware Projection with Arbitary Graphs